A water tank can be filled in 24 hours by an inlet pipe and emptied in 30 hours by an outlet pipe. How long will it take to fill the tank if both pipes are left open?
Answer by addingup(3677)

1/24-1/30 = 1/x Multiply each term by LCD 120x
(1*120x)/24-(1*120x)/30 = (1*120x)/x
5x-4x = 120
x = 120 hours